# Don’t be lazy: CompleteP enables compute-efficient deep transformers

### v1

@@ Appendix B Limitations
Our theoretical analysis considers scaling width and depth at fixed token count, which limits its the direct applicability to the fixed TPP compute optimal regime. However, empirically the scaling predictions derived in the fixed token setting have great .
@@ E.4 Adam $\epsilon$ Parameter
These rules will ensure the correct scale updates in each layer. We summarize this global $\epsilon$ strategy in Table 1 and empirically validate it in Appendix F.
@@ I.5 Figure 6 details
We then perform a single gradient step only with respect to $\mathbf{W}_{(2)}^{\ell}$ for a fixed layer $\ell$, and report the median on the left hand side of the laziness equation above for varying depths across $50$ seeds per run. We compute $\Delta_{\theta}\mathbf{h}_{\ell}^{lin,\theta}$ explicitly by differentiating the block with respect to $\mathbf{W}_{(2)}^{\ell}$ and evaluate at the initial parameter values, i.e. we compute:
@@
where $\bm{\theta}=\mathbf{W}_{(2)}^{\ell}$. We use $\eta_{0}=0.0001$ and train on MNIST [106]. We provide the code to reproduce this plot at https://github.com/EleutherAI/nanoGPT-mup/blob/completeP/mup_examples/convergence_to_linearized_model.ipynb.

### v2

@@ Appendix B Limitations
Our theoretical analysis considers scaling width and depth at fixed token count, which limits the direct applicability to the fixed TPP compute-optimal regime. However, the scaling predictions derived in the fixed token setting have predicted success in the fixed TPP compute-optimal regime.
@@ E.4 Adam $\epsilon$ Parameter
These rules will ensure the correct scale updates in each layer.
@@ I.5 Figure 6 details
We then perform a single gradient step only with respect to $\mathbf{W}_{(2)}^{\ell}$ for a fixed layer $\ell$, and calculate the left hand side of the laziness equation above for varying depths across $50$ seeds per run. We report the median with lines and shade the region between the first and third quartiles. We compute $\Delta_{\theta}\mathbf{h}_{\ell}^{lin,\theta}$ explicitly by differentiating the block with respect to $\mathbf{W}_{(2)}^{\ell}$ and evaluate at the initial parameter values, i.e. we compute:
@@
where $\bm{\theta}=\mathbf{W}_{(2)}^{\ell}$. We use $\eta_{0}=0.0001$ and train on MNIST [107]. We provide the code to reproduce this plot at https://github.com/EleutherAI/nanoGPT-mup/blob/completep/completep_examples/relative_distance_from_linearization.ipynb.